Categories

keeping the bees busy

Today marked the beginning of our third week of the field season. The team began the day with phenology measurements in the remnants. Today we measured 509 flowering Echinacea heads! This is roughly 1/4 to 1/3 of the heads we are monitoring. We expect flowering will peak next week. Stay tuned for more information! During the afternoon, the team worked on independent projects and continued assessing Hesperostipa reproduction in remnants. Emma and Amy measured a plot at landfill east with over 100 culms… oofta!

We are staying busy, and so are the bees!

Facebooktwitterpinterestmail

Leave a Reply

You can use these HTML tags

<a href="" title=""> <abbr title=""> <acronym title=""> <b> <blockquote cite=""> <cite> <code> <del datetime=""> <em> <i> <q cite=""> <s> <strike> <strong>